# FLAF: Focal Line and Feature-constrained Active View Planning for Visual Teach and Repeat

### v1

@@ IV-B Tracking Failure Avoidance Validation
As shown in Table I and Fig. 5, the passive VT&R achieves a stable and accurate performance on the first two paths but fails in the teach phase on path three. The few low-texture regions on path one and two are avoided by a considerable human guide. Our active VT&R system with active view planning succeeds in the teach phase on three paths.
Fig. 6 demonstrates the situations where low-texture regions are overcome by our active VT&R with an active camera that automatically captures informative regions to maintain a stable localization. At position one, the active camera automatically looks at the poster in the upper left to avoid the white wall. At position two, the active camera looks up at the ceiling to maintain the localization relying on the square lamps. The robot looks toward the upper right at position three to focus on the logo while passing through a low-texture corner. Finally, at position four, the robot looks toward the upper left at the scroll on the door for abundant features.
IV-C Active View Planning Method Evaluation
Active SLAM with different observation models typically succeeds in the teach phase on all paths. Because they mostly control the robot to see the partial map just built. This is achieved by looking at the directions that have more mapped features. However, existing methods such as UDVP[16] ignore the identifiability of the points by feature extractor and matcher. In the mapping phase active SLAM has no choice but to see the just mapped points because there are no more feature points triangulated. However, in the repeat phase, some of the map points (3D coordinates) can’t be identified by VSLAM from a large view angle compared with the view direction in which they were triangulated.
Our FLAF-constrained active view planning outperforms the UDVP model in repeating a complete path because of our consideration of the affine change of the feature points. The map point’s normal line ($\mathbf{n}_{p}$) shown in Fig. 3 limits the orientation of the active camera in the view angle-invariance range of the feature. Table I shows that our method repeats a more complete trajectory on all paths, especially in some challenging cases. Although on path three UDVP achieves more accurate path tracking, the trajectory errors are all negligible for VT&R. The trajectories shown in Fig. 5 also verify that the repeat consistency with different methods are all pretty good, which means the repeat trajectories always fit with the taught one. Part of the errors are caused by the time-wise point-to-point comparison. The timestamps are difficult to align. We just modified them to be uniformly consistent at the same time duration. Initially, we implement our FLAF model by counting the map points in the FoV of a PTU sample that accords with FLAF, which is more accurate for the higher speed. Only the final version of our FLAF-constrained active VT&R successfully finishes all three paths.
Figure 7: Number of matched points in local map tracking with respect to time and the failure case of UDVP method on path 2 shown in feature map. Green square represents current pose of the camera.
@@ IV-D Map Points Association Validation
In[12], the authors draw a point graph to show the relationship between the probability of the tracking failure and the number of observed map points. This work[12] indicates that the probability of the tracking failure approaches zero if the observed map points are more than a threshold. Our work further indicates that the tracking failure is caused by failed local map tracking. We use the analyzing method proposed in[12] to analyze the state of map point association in the repeat phase. As shown in Fig. 7, we recorded the quantity of matched points in local map tracking and drew the line graph of two methods for all three paths.
On path one, the active camera-based VSLAM with FLAF matches fewer points than it is by UDVP initially. However, the VSLAM with UDVP fails to track the path after a reduction of matched points. On the contrary, the VSLAM with our proposed FLAF-based method initially tracks fewer points of the local map but maintains a stable localization and increasing number of matches. On path three, the active visual repeat with these two view planning methods tracks a similar quantity of map points, but our method recovers the localization after a challenging decline where the UDVP-based method fails.
On path three, a similar case happens that the UDVP method loses the localization after a reduction of matched map points. In the same challenging location, our FLAF-based method finishes repeating the path by adjusting the active camera to the direction with more identifiable map points. Fig. 7 also shows that the UDVP method controls the active camera to see the direction with more local map points without considering the feature identifiability of these points. It means although these points are in the FoV of the camera, they can not be identified and matched by the feature extractor and matcher because of the ignorance of the affine change of the points.
@@ V Conclusion
V-A Conclusions
In this research, we present a novel active view planning method for VT&R to solve the tracking failure caused by low-texture regions. Our passive VT&R system is built on feature-based VSLAM and faces the main problem of tracking failure caused by low-texture regions. Existing feature-based active VT&R systems use a gimbal-stabilized camera to improve the stability. The gimbal is typically manually steered in the teach phase instead of using active view planning. Previous works propose some active view planning methods for VSLAM. However, these works haven’t demonstrated how to use these methods to help VT&R with an active camera.
We aim to use an active camera with active view planning to build up an active VSLAM system that can support our active VT&R system. After our test, a recent active view planning method for VSLAM succeeds in the teach phase but easily fails in the repeat phase. This method controls the camera toward directions with more points close to the camera. However, it fails to track the local map in cases when the closer points are not identifiable by the ORB feature extractor and matcher. Therefore, we design a focal line and feature (FLAF)-constrained active view planning method taking into account the affine change of feature points.
According to our experimental results, passive VT&R fails with the low-texture regions. Active camera-based VSLAM with an existing view planning method easily fails in the repeat phase because it doesn’t account for view angle-dependent affine change of the features. With our proposed FLAF model, the novel active view planning method successfully helps our VT&R system finish all three paths. Our experimental results also demonstrate the process of overcoming the low-texture regions with an active camera. Finally, we analyze the reason for tracking failure through three graphs that show the relationship between the quantity of matched points and time.
Our most important contribution is the demonstration of active camera-based VT&R. We focus on solving the problem of tracking failure in VT&R caused by low-texture regions. We will address the dilemma between exploitation of the existing map and exploration of the unknown environment in the future.

### v2

@@ IV-B Tracking Failure Avoidance Validation
As in Table I and Fig. 5, the passive VT&R achieves a stable and accurate performance on the first two paths but fails in the teaching phase on paths 3 and 4. The few low-texture regions on path one and two are avoided by a considerable human guide. Our active VT&R system with three view planner succeeds in the teaching phase on all 4 paths.
Fig. 6 demonstrates the situations where low-texture regions are overcome by our active VT&R with an active camera that automatically captures informative regions to maintain a stable localization. At position 1, the active camera automatically looks at the poster in the upper left to avoid the white wall. At position 2, the active camera looks up at the ceiling to maintain the localization relying on the square lamps. The robot looks toward the upper right at position 3 to focus on the logo while passing through a low-texture corner. Finally, at position 4, the robot looks toward the upper left at the scroll on the door for abundant features.
IV-C Active View Planning Method Comparison
Active SLAM with different observation models typically succeeds in the teaching phase on all 4 paths. The active camera-based VSLAM has no choice but to see the just mapped points while there are no more feature points triangulated. However, in the repeat phase, existing methods such as UDVP[16] ignore the identifiability of the points by feature detector. Some of the map points (3D coordinates) cannot be identified by VSLAM from a large view angle compared with the view direction in which they were triangulated.
Our FLAF-constrained active view planning outperforms the UDVP method in repeating complete paths because it accounts for the affine change of feature points. The normal line of the map point ($\mathbf{n}_{p}$), shown in Fig. 3, limits the orientation of the active camera in the view angle-invariant range of the feature. As seen in Table I, our method repeats more complete trajectories across all paths, especially in challenging scenarios. Although UDVP achieves more accurate path tracking on path 3, the trajectory errors on all paths by all view planners are negligible for VT&R. Part of the AP-RMSE arises from time-wise point-to-point comparisons, which are challenging to execute precisely. To address this, we modified the timestamps to be uniformly consistent over the same time duration. The trajectories shown in Fig. 5 further confirm that the repeat consistencies with different methods are all pretty good, as the repeated trajectories align well with the taught ones.
Figure 7: Number of matched points in local map tracking with respect to time. The bottom shows the failure case of the UDVP method on path 2 shown in a feature map. The green square represents the current pose of the camera.
@@ IV-D Map Points Association Validation
In[12], the authors present a line graph illustrating the relationship between the probability of the tracking failure and the number of observed feature points. Their results[12] indicate that the likelihood of the tracking failure approaches zero when the number of associated map points exceeds a certain threshold. From the perspective of our work, the specific failure is caused by choosing the wrong orientation of the active camera relative to the local map. To further investigate, we applied the analysis method proposed by [12] to examine the state of map point association during the repeating phase.
As shown in Fig. 7, we recorded the number of inlier points successfully matched during local map tracking and plotted line graphs comparing the performance of the two methods across all test paths. On paths 1 and 2, the active camera-based VSLAM with FLAF initially matched fewer points than the UDVP-based system. However, after an initial reduction of matched points, the UDVP-based method fails to maintain tracking, while the FLAF-based method continues to provide stable localization and an increasing number of matches. This phenomenon suggests that, once the inlier number of the tracked map points reaches the threshold, the sheer number of map points becomes less critical for stable localization. The bottom of Fig. 7 also illustrates how the UDVP method directs the active camera toward regions with more local map points, without accounting for the feature identifiability. Although many points may fall within the camera’s FoV, they may not be recognized or matched by the feature extractor due to the ignorance of the affine changes.
On paths 3 and 4, the active visual repeat with both the FLAF and UDVP view planners tracks a similar quantity of map points. However, our FLAF-based method successfully recovers localization after a challenging decline in tracked points, where the UDVP-based approach fails. Even in cases of temporary tracking loss, the active camera controlled by our view planner was able to orient itself appropriately by executing the instruction before tracking loss, allowing the VSLAM system to recover localization through place recognition and the PnP algorithm.
@@ V Conclusion
In this research, we present a novel active view planning method for VT&R that addresses the tracking failure caused by the low-texture regions and demonstrates the whole active VT&R. Our experimental results show that our active VT&R successfully overcame the specific failure of passive VT&R and our proposed FALF-constrained active view planning outperforms existing view planners in completion and success rate of VT&R.
A previous active camera-based VT&R manually steered a gimbal-stabilized camera instead of rotating actively for failure avoidance. Existing active camera-based VSLAM methods did not demonstrate how to integrate their View planners into visual navigation systems. During our tests, the VT&R systems built on these active camera-based VSLAM systems easily failed in the repeat phase without considering the feature-identifiability of the map points. Our proposed focal line and feature (FLAF)-constrained active view planning successfully addressed these failures by considering the view angle difference between the current viewpoint and those at which the map points were triangulated. With our view planner, the active VT&R system successfully finishes all four paths at the highest completion and success rate. Additionally, our point-line plots indicate that the quality of the map points is more important than quantity for stable localization.
However, for each execution of view planning, 900 samples of PTU angles were scored according to thousands of map points in the local map, resulting in a high computational overhead. Despite using OpenMP to speed up the view planners, they operated at less than 5 FPS, which hindered the performance of VSLAM and reduced the success rate of VT&R by preventing accurate path reconstruction. In future work, we will address this limitation by parallelizing the view planning module with the VSLAM system to improve processing speed and overall system performance.